OpenStudy (anonymous):

whats the square root of a^2/b^5

OpenStudy (mertsj):

\[\sqrt{\frac{a^2}{b^5}}=\frac{\sqrt{a^2}}{\sqrt{b^4b}}\]

OpenStudy (mertsj):

You can finish the simplification.

OpenStudy (anonymous):

\[\large \sqrt{a \over b}={\sqrt{a} \over \sqrt{b}}={a^{1/2} \over b^{1/2}}\]\[\large (a^x)^y=a^{xy}\]

OpenStudy (anonymous):

If you let me know what you get I'll check it for you. Use the rules I outlined above

OpenStudy (anonymous):

@ChmE I'm still stuck on this problem because its not using numbers, I know how to do something like